# Conghui He1,3‡, Weijia Li2

### v2

@@ Introduction
Figure 1: Comparison between UrBench and previous works. (1) UrBench contains both region-level and role-level questions, while previous benchmarks generally focus on region-level questions. (2) In addition to single-view questions in satellite view or street view, UrBench also incorporates cross-view questions, in which each question contains images of the same scenario from different views. (3) It evaluates LMMs on a comprehensive range of 14 diverse tasks in 4 evaluation dimensions.
@@
Question Test Val
Total 10483 1117
View Type cross 4208 434
sat 2618 268
mono-str 1694 195
pano-str 1735 196
multi-str 228 24
Visual Prompt with 4227 446
without 6256 671
Answer Type MC 8920 964
open 1563 153
Images per Q 1 image 5522 606
2 images 3485 357
4 images 753 77
5 images 723 77
@@
Building on the strengths of Large Language Models (LLMs) (Brown et al. 2020; Chowdhery et al. 2023; Touvron et al. 2023) in complex language reasoning and understanding, Large Multimodal Models (LMMs) can process inputs from multiple modalities and accomplish sophisticated visual reasoning and understanding tasks (Yin et al. 2023; Bai et al. 2024). The rapid growth of LMMs has given rise to both closed-source models like GPT-4o (Achiam et al. 2023) and Gemini (Reid et al. 2024), as well as open-source models such as the InternVL series (Chen et al. 2024), LLaVA series (Liu et al. 2024b), and VILA series (Lin et al. 2024), all of which have demonstrated significant potential for various application tasks (Cui et al. 2024; Xiao et al. 2024).
@@
However, these works primarily focus on LLMs alone or in combination with CLIP, with limited analysis of LMMs. Additionally, they concentrate on a narrow range of single-view urban tasks that do not demand advanced visual reasoning or multi-view comprehension. This leaves a significant gap in the comprehensive evaluation of LMMs’ performance on urban-related tasks. Our objective is to assess LMMs in a wider variety of urban tasks, broadening the scope to include more extensive view coverage and fully harnessing their potential in urban scenarios.
@@
Geo-Localization. This dimension contains role-level tasks widely used in remote-sensing (Zhu, Yang, and Chen 2021), which requires LMMs to predict geographical coordinates and directions given images (Task 1-4 in Fig.4). For example, in Image Retrieval (IR), we query LMMs with satellite or street view images to retrieve their corresponding counterpart, while City Retrieval (CR) tasks LMMs to predict the name of a city given its satellite or street view images. Both Orientation (OR) and Camera Localization (CR) require LMMs to pinpoint directions given references in cross-view images.
@@
Benchmark Curation
Data Pre-processing. In this stage, we process our collected raw image data to produce annotations for later stages. For cross-view tasks that require object matching, we develop a cross-view detection-matching method to obtain paired-up instance level annotations. Specifically, We first use a pretrained Grounding DINO (Liu et al. 2023a) to obtain bounding box annotations for street-view images. Since bounding box annotations for satellite-view images are already obtained through OSM, we apply ray tracing to map the street-view boxes to the satellite view. Next, we calculate the IoUs between the mapped and original satellite-view boxes and select pairs with IoUs over 0.5 as cross-view matches. We then employ human checking to further ensure the quality of our matching algorithm. This way, we effectively align bounding boxes from different views at an instance level. Additionally, we construct a comprehensive annotation database by unifying annotations from different datasets, which facilitates subsequent data generation and quality control.

### v3

# UrBench: A Comprehensive Benchmark for Evaluating Large Multimodal Models in Multi-View Urban Scenarios

@@ Introduction
Figure 1: Comparison between UrBench and previous works. (1) UrBench contains both region-level and role-level questions, while previous benchmarks generally focus on region-level questions. (2) In addition to single-view questions in satellite or street view, UrBench also incorporates cross-view questions. (3) It evaluates LMMs on a comprehensive range of 14 diverse tasks in 4 evaluation dimensions.
@@
Building on the strengths of Large Language Models (LLMs) (Brown et al. 2020; Touvron et al. 2023) in complex language reasoning and understanding, Large Multimodal Models (LMMs) can process inputs from multiple modalities and accomplish sophisticated visual reasoning and understanding tasks (Yin et al. 2023; Bai et al. 2024). The rapid growth of LMMs has given rise to both closed-source models like GPT-4o (Achiam et al. 2023) and Gemini (Reid et al. 2024), as well as open-source models such as LLaVA (Liu et al. 2024b), and VILA (Lin et al. 2024), all of which have demonstrated significant potential for various application tasks (Cui et al. 2024; Xiao et al. 2024).
@@
Geo-Localization. This dimension contains role-level tasks widely used in remote-sensing (Zhu, Yang, and Chen 2021), which requires LMMs to predict geographical coordinates and directions given images (Task 1-4 in Fig.4). For example, in Image Retrieval (IR), we query LMMs with satellite or street view images to retrieve their corresponding counterpart, while City Retrieval (CR) tasks LMMs to predict the name of a city given its satellite or street view images. Orientation (OR) and Camera Localization (CR) require LMMs to pinpoint directions given cross-view information.
@@
Benchmark Curation
Data Pre-processing. In this stage, we process our collected raw image data to produce annotations for later stages. For cross-view tasks that require object matching, we develop a cross-view detection-matching method to obtain paired-up instance level annotations. Specifically, We first use a pretrained Grounding DINO (Liu et al. 2023a) to obtain bounding box annotations for street-view images. Since bounding box annotations for satellite-view images are already obtained through OSM, we apply ray tracing to map the street-view boxes to the satellite view. Next, we calculate the IoUs between the mapped and original satellite-view boxes and select pairs with IoUs over 0.5 as cross-view matches. We then employ human checking to further ensure the quality of our matching algorithm. This way, we effectively align bounding boxes from different views at an instance level. Additionally, we construct a comprehensive annotation database by unifying annotations from different datasets, facilitating data generation and quality control.
@@ Conclusion
Acknowledgments
This project was funded by National Natural Science Foundation of China (Grant No. 42201358) and Shanghai AI Laboratory.